by TEI TEI

Partner Lead: TEI

10/31/13 – Issued for bidding on 10/31/13.

01/24/14 – Contracts set to Bert Watts, waiting for approval from prosecutor to schedule pre-construction meeting.

02/25/14 – Contract in Phasor’s court to be signed and returned and sent to purchasing for purchase order. Tentative pre-construction date is 03/19/14 at 2:00 PM

03/26/14 – Project went to BOCC and was approved. Waiting for contracts and PO to schedule pre-construction meeting.

04/22/14 – Project preconstruction meeting scheduled for Monday 4/28/14 at 2PM, meet in the south tower lobby to discuss schedule, background checks, etc.

05/28/14 – TEI approved all submittals on 05/27/14 and Bert Watts issued Notice to Proceed on 05/28/14. Phasor to provide schedule and notice of all BGC from ICS.

06/25/14 – ICS has all of the head end equipment at their ship and is in the process of programming all of the equipment. A project kickoff with Bldg personnel, Sheriff’s personnel, Facilities, TEI and Phasor to be scheduled mid next week. Anticipated start date is on or around July 14, 2014.

07/30/14 – Phasor mobilized to Justice Center for work on this project. Currently installing all of the head end racks in each of the LC’s. Upon completion of install of the racks, and bolting in place, then Phasor to route 120 power to each of the racks. construction meeting scheduled for Aug 6, 2014 at 1:00 PM.

8/25/14 – All racks are installed in place in all of the LC’s. Phasor de-mobilized for approximately 2 weeks waiting on the delivery of the speakers for replacement in the pods. ICS received the first shipment (approximately 200 units) the week of 08/18, and the second shipment of approx 200 units has shipped and is scheduled to be delivered the week of 08/25. The order is scheduled to be completely fulfilled no later than 09/05/14. Phasor has re-mobilized on 08/25/14 to begin the replacement of the speakers, starting with the problematic unit in George 51.

09/24/14 – In the North Tower we have 289 new installed and approximately 140 old speakers left until we determine which we are going to use the Lexan plates on. In the South building, they have went through Lincoln 52 and were able to install 51 new and 5 existing were left in. They are currently working on Lincoln 51.

10/28/14 – The north tower is 100% complete with installation. The south tower is 100% complete with installation of systems on 3,4 and 5. ICS has placed order for balance of speakers that were not included on original order (receipt of speakers is approximately 10 weeks out – first of 2015). Phasor to pull off project after 10/29/14 punchlist with ThermalTech until such a time that speakers are on site.Phasor stated that the amount of remaining lexan plates should be adequate for the remaining scope of the project.

01/27/15 – The south tower is 100% complete with installation. TEI completed punchlist on 1/26/15 of the south tower. Phasor and ICS to complete punchlist and provide closeout documentation.

02/24/15 – Phasor submitted O&M manuals and as-built drawings to ThermalTech for review and submitting to the county. O&M’s turned over at the 02/26/15 meeting.

03/25/15 – project complete and can be archived.